OpenStudy (anonymous):

Help urgent!!!! Approximate the arclength of y= e^(2x) from x = 0 to x= 3 using Trapezoidal Rule (n=6).

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \int_{a}^{b} f(x)~dx~~~~{\normalsize~~using~~{\rm n}~~number~~of~~shapes} }\) \(\large\color{black}{ \displaystyle \int_{a}^{b} f(x)~dx=\normalsize ~\frac{1}{2}\cdot \frac{b-a}{n}\cdot \left(y_0+2y_1+2y_2+2y_3+...+2y_{n-1}+y_n\right) }\)